During 2013, the country of Economic had a real GDP of $115 billion and the population was 0.9 billion. In 2012, real GDP was 105 billion and the population was 0.85 billion. Economics growth rate of real GDP per person is
1
Expert's answer
2020-04-29T09:20:49-0400

2013=GDP of $115 billion and population=0.9billion

2012==GDP of $105 billion and population=0.85 billion

Economics growth rate of real GDP, Egr per person is;

"Egr=\frac{\text{change in real GDP}}{\text{change in population}}"

Egr=(10/0.05)= 200 GDP/population